# On the interaction of dilatancy and friction in the behavior of fluid–saturated sheared granular materials: a coupled Computational Fluid Dynamics–Discrete Element Method study

### v1

@@ IV Spatial and temporal distribution of friction, pore pressure, and volume change
Figure 7: Local rheology for subaerial and subaqueous models under dense and loose conditions. The color gradient from dark to light represents the temporal evolution from the onset of failure to the final stage of deformation. (a) Variation of the friction coefficient with inertial number ($I_{n}$) and deviatoric strain rate for subaerial models, showing rate-strengthening behavior. (b) Variation of the friction coefficient with viscous number ($I_{v}$) and deviatoric strain rate for subaqueous models, also indicating rate strengthening, with higher frictional values compared to subaerial cases. (c) Variation of porosity with inertial number ($I_{n}$) and deviatoric strain rate for subaerial models. (d) Variation of porosity with viscous number ($I_{v}$) and deviatoric strain rate for subaqueous models.
@@ V.1 Dry condition
Frictional behavior in subaerial granular systems is strongly influenced by strain rate and the inertial number ($I_{n}$), reflecting the dominant role of dry interparticle interactions. Dense subaerial materials exhibit a modest but consistent increase in friction with rising strain rate (Fig. 7(a)), characteristic of a rate-strengthening response, with the color gradient representing the temporal evolution from the onset of failure (dark) to the final stage (light) of deformation. This behavior stems from tightly packed grain assemblies that promote the development of persistent force chains and robust interparticle contacts, which enhance resistance to shear deformation.
@@
In contrast, loose subaerial materials display lower friction at low strain rates due to limited and transient grain contacts. As strain rate increases, friction rises more markedly, driven by progressive particle rearrangement, localized compaction, and the formation of new contact networks throughout the deformation process (Fig. 7(a)), with the color gradient representing the temporal evolution from the onset of failure (dark) to the final stage (light) of deformation. This evolving frictional response reflects the gradual restructuring of the granular framework in response to increased shear input.
@@
When friction is plotted against the inertial number, both dense and loose materials show an increasing trend (Fig. 7(a)). However, the increase in loose materials is more gradual and becomes more pronounced only at higher $I_{n}$ values. These trends underscore the importance of packing density in governing strain localization, contact mechanics, and frictional evolution within dry granular systems.
@@
While porosity in granular media is generally expected to vary with strain rate due to the competing effects of compaction and dilation, it can remain relatively stable over certain regimes. This constancy arises when the system achieves a balance between structural compression and dilation, or when strain rates are too low to significantly disrupt the grain arrangement. In the dense model, porosity remains nearly constant for inertial numbers below $10^{-1}$ (Fig. 7(c)). Parez et al. (2021) reported a similar trend, observing porosity stability across $I_{n}$ values ranging from $10^{-6}$ to $10^{-2}$, consistent with our findings. Likewise, Amarsid et al. (2024) found that porosity remains largely unaffected at low inertial numbers, further corroborating these observations for both dense and loose conditions.
V.2 Fluid saturated condition
Subaqueous granular rheology is governed by the intricate interplay between solid grains and the surrounding fluid. In these systems, friction emerges from both interparticle contact forces and hydrodynamic interactions, introducing greater complexity compared to dry granular materials.
@@ V.2 Fluid saturated condition
Loose subaqueous granular materials exhibit a strain-rate-dependent response characterized by two distinct regimes (Fig. 7(b)). At the onset of shearing, increasing strain rates lead to pore-pressure buildup, which enhances friction through fluid drag and the development of reinforced contact networks. Conversely, at lower strain rates, typically observed in the later stages of shearing, fluid lubrication dominates. This lubrication reduces friction by facilitating easier particle rearrangement. The observed rate-strengthening behavior underscores the importance of fluid-mediated forces in controlling deformation within loosely packed systems.
@@
Dense subaqueous materials respond differently, exhibiting comparatively lower friction with increasing strain rate (Fig. 7(b)), primarily due to the development of negative pore pressure at the onset of shear. As strain rate continues to rise, fluid drag becomes increasingly influential, and pore-pressure effects become more evident, reflecting the evolving interaction between the solid framework and the interstitial fluid. When friction is plotted against the viscous number ($I_{v}$), dense materials tend to stabilize at higher $I_{v}$ values, showing reduced frictional variability. This plateau suggests that viscous forces play a moderating role in the frictional response under steady fluid–particle interactions. In contrast, loose materials display a more pronounced increase in friction with rising $I_{v}$, driven by elevated resistance due to intensified fluid drag and pore-pressure buildup.
@@
Although porosity in fluid-saturated granular materials might be expected to vary with strain rate, observations show that porosity can remain nearly constant over a moderate range of strain rates (Fig. 7(d)). This behavior arises when the system reaches a dynamic balance between compaction and dilation, or when the imposed strain rates are too low to significantly alter the microstructure. Amarsid et al. (2024) similarly found that porosity, governed by the viscous number, remains largely unaffected at lower $I_{v}$ values, consistent with the trends identified here.
@@
A direct comparison of dry and saturated systems highlights the critical role of fluid presence and initial structure in governing shear response: while subaerial systems rely on interparticle contact mechanics, subaqueous systems are shaped by a complex feedback between fluid drag, pore pressure, and microstructural evolution.
@@ VI Analytical solution for excess pore pressure using poromechanics
During breaching, a dynamic equilibrium emerges between pore pressure dissipation and continuous dilation You, Flemings, and Mohrig (2012b); Alhaddad et al. (2024). As pore pressure dissipates, the effective stress within the slope deposit increases, stabilizing the sediment. However, for retrogressive failure to persist, this dissipation must be sustained by ongoing dilation. The resulting feedback mechanism, where slope failure induces dilative strain that regenerates negative excess pore pressure, helps maintain the breaching process. Importantly, this dynamic decouples the steady-state pore pressure distribution from the deposit’s intrinsic permeability, revealing an underexplored aspect of sediment failure behavior You, Flemings, and Mohrig (2012b); Alhaddad et al. (2024).
To formalize this equilibrium, the analytical solution developed by You, Flemings, and Mohrig (2012b) describes the pore pressure evolution coupled with dilation during the breaching process. This formulation predicts the steady-state negative pore pressure field, showing that the system can self-regulate pore pressure independently of permeability once dilation-driven feedback is active.
Comparisons with large-scale laboratory experiments and numerical models confirm this feedback mechanism’s role in sustaining breaching and highlight deviations during the transient initiation phase Alhaddad et al. (2024, 2023). The same analytical formulation is applied in this study to evaluate whether the CFD–DEM simulation results for excess pore pressure in both dense and loose subaqueous granular models follow the theoretical predictions, particularly at steady state. By extracting simulation data at various locations and times, we directly compare the evolution and final distribution of pore pressure against the analytical model.
Furthermore, recent hydromechanical models of subduction zone sediments Nikolinakou et al. (2023) demonstrate that shear-induced overpressures, in addition to changes in mean stress, significantly control pore pressure evolution. Although applied in a tectonic context, these findings conceptually reinforce that dilative strain and shear play a dominant role in transient and steady-state pore pressure fields, complementing observations from breaching processes.
@@
$\hat{u}^{*}=\frac{\delta\beta s_{0}}{1-\delta}\left(e^{-\eta\hat{x}}-e^{-\frac{v}{C_{v}}\hat{x}}\right)$ (18)
Eq. 17 describes pore pressure dissipation through Darcy flow with a moving boundary (first two terms) and continuous dilation effects (third term). The coefficient of consolidation, $C_{v}$, combines the effects of permeability, viscosity, and compressibility. $\hat{u}$ represents the excess pore pressure, $\hat{\sigma}_{3}$ is the least principal stress (assumed horizontal), and $C_{v}$ is the coefficient of consolidation. The boundary conditions applied in this analytical formulation are critical to reproducing the steady-state pore pressure field. In the Lagrangian reference frame, where the coordinate system moves with the retreating failure front, hydrostatic pore pressure is assumed at both the upstream (breaching front) and downstream (far-field) boundaries You, Flemings, and Mohrig (2014b). Additionally, no-flow conditions are imposed at the base and at the right boundary of the computational domain, assuming symmetry and minimal lateral drainage. These boundary constraints ensure that pore pressure evolves solely due to dilation-induced stress changes and fluid migration within the domain. The steady-state solution derived under these conditions shows that the excess pore pressure depends only on the dilation potential, effective stress, and retreat velocity, but not on permeability once equilibrium is reached. This insight reveals that boundary-controlled feedback mechanisms can dominate the evolution of excess pore pressure during breaching, consistent with both pore pressure measurements and model predictions You, Flemings, and Mohrig (2014b). Its solution, given by Eq. 18, represents $\hat{u}^{*}$, the excess pore pressure in the Lagrangian reference frame. The parameter $\beta$ denotes the relative dilation strength, defined as the ratio between differential stress-induced strain and mean effective stress-induced strain. $s_{0}$ represents the far-field horizontal stress. The coefficient of consolidation is defined as $C_{v}=\frac{k}{\mu E}$, where $k$ is the permeability, $\mu$ is the dynamic viscosity of the pore fluid, and $E$ is a modulus related to volumetric strain. The coordinate $\hat{x}$ is defined in the Lagrangian reference frame as $\hat{x}=x-vt$, where $v$ is the velocity of the retreating breaching front (Eq. 24). $\delta$ characterizes the ratio between $v$ and the product of the effective stress decay rate $\eta$ and the coefficient of consolidation $C_{v}$. The parameter $\eta$ is a constant with units of inverse length, affecting the spatial distribution of excess pore pressure.
@@
Figure 9: Evolution of excess pore pressure response during breaching and comparison with the analytical solution: (a) Monitored points A, B, C, D, E, and F for the subaqueous dense model; (d) monitored points G, H, I, J, K, and L for the subaqueous loose model. (b) In the dense model, wall removal results in an abrupt drop in pore pressure, followed by a gradual rise as the system stabilizes. (e) In contrast, the loose model exhibits an abrupt increase in pore pressure after wall removal, followed by a gradual decline. (c) and (f) Excess pore pressure at time intervals of 0.01 s, 0.1 s, 0.4 s, and 0.8 s (with dashed lines indicating the onset of breaching) is plotted against the distance from the breaching front using Lagrangian coordinates for the dense (point D) and loose (point J) models, respectively. The solid line represents the analytical solution. The comparison shows good agreement at steady state for both models: in the dense model, pore pressure begins at a negative value, increases, and stabilizes; in the loose model, pore pressure starts high, decreases, and reaches a stable value. Transient phases, however, show less agreement with the analytical model, highlighting the complexity of early-stage fluid–particle interactions.
@@
Eq. 20 defines permeability $k$ as a function of fluid properties and flow path geometry, where $v_{\text{avg}}$ is the average velocity, $L$ is the flow path length, and $\Delta P$ is the pressure difference. The calculated permeability values are $k=1.75\times 10^{-7}\,\mathrm{m}^{2}$ for the loose model and $k=0.25\times 10^{-7}\,\mathrm{m}^{2}$ for the dense model. The evolution of permeability over time at selected points, derived using Darcy’s law, is presented in Supplementary Fig. S1(a). In Eq. 21, $\beta$ is the dimensionless dilation strength, where $\beta>\frac{1}{2}$ indicates dilative material. Eq. 23 describes the least principal stress, which decays exponentially with distance from the breaching front, explaining localized pore pressure changes during continuous dilation. During breaching, the volumetric strain rate equals the water flux into the deposit, and strain is proportional to the minimum pore pressure. The parameter $s_{0}$ represents the least principal stress as $x\to\infty$, and $\eta$ defines the rate of stress decay.

### v2

@@ IV Spatial and temporal distribution of friction, pore pressure, and volume change
Figure 10: Local rheology for subaerial and subaqueous under dense and loose conditions. The color gradient from dark to light represents the temporal evolution from the onset of failure to the final stage of deformation. (a) Variation of the friction coefficient with inertial number ($I_{n}$) and deviatoric strain rate for subaerial models, showing rate-strengthening behavior. (b) Variation of the friction coefficient with viscous number ($I_{v}$) and deviatoric strain rate for subaqueous models, also indicating rate strengthening, with higher frictional values compared to subaerial cases. (c) Variation of porosity with inertial number ($I_{n}$) and deviatoric strain rate for subaerial models. (d) Variation of porosity with viscous number ($I_{v}$) and deviatoric strain rate for subaqueous models.
The results above pertain to non-cohesive, quartz-sand–like grains, focusing on the coupling between friction, pore pressure, and tendency for volume-change. Many natural submerged sediments include clayey and silt-sized particles that are known to demonstrate cohesive interactions. The presence of such cohesive interactions is expected to suppress dilation, raise peak friction (Herminghaus2005) at the onset of deformation of materials, shorten runout (Zhu2022), and subaqueously reduce effective permeability and slow pore-pressure dissipation (Vowinckel2019). Cohesion also tends to narrow shear bands in dense flows Zhao2023. Grain-resolved studies report reduced mobility and altered excess-pressure evolution in cohesive beds Vowinckel2019; Zhu2022. A cohesive extension of our study is planned as future work, focusing on the systematic exploration and comparison of failure dynamics with cohesive-bed benchmarks (Vowinckel2019).
@@ V Granular rheology in shear zone
Figure 11: Unified rheology across subaerial (inertial number $I_{n}$) and subaqueous (viscous number $I_{v}$) conditions using the combined number $K=I_{v}+\alpha I_{n}^{2}$. Friction coefficient $\mu$ as a function of $K$; the curve follows Eq. 19. Fitted parameters are reported in the text.
Figure 12: Friction–packing trends across subaerial and subaqueous flows. Symbols show friction $\mu$ versus packing fraction $\phi_{s}$, colored by $x=\log_{10}(I_{n}/I_{v})$, where $I_{n}$ and $I_{v}$ are the inertial and viscous numbers, respectively.
Figure 13: (a) Friction coefficient $\mu$ versus $x=\log_{10}(I_{n}/I_{v})$ for subaerial and subaqueous flows. (b) Packing fraction $\phi_{s}$ versus $x$. The vertical dashed line marks the crossover $I_{n}=I_{v}$ ($x=0$). Shading indicates regimes: inertial ($x>0$, light gray) and viscous ($x<0$, light blue).
@@ V.1 Dry condition
Frictional behavior in subaerial granular systems is strongly influenced by strain rate and the inertial number ($I_{n}$), reflecting the dominant role of dry interparticle interactions. Dense subaerial materials exhibit a modest but consistent increase in friction with rising strain rate (Fig. 10(a)), characteristic of a rate-strengthening response, with the color gradient representing the temporal evolution from the onset of failure (dark) to the final stage (light) of deformation. This behavior stems from tightly packed grain assemblies that promote the formation of persistent force chains and robust interparticle contacts, thereby enhancing resistance to shear deformation.
@@
In contrast, loosely packed subaerial materials exhibit lower friction at low strain rates because grain contacts are sparse and short-lived. As the strain rate increases, friction rises more sharply, driven by progressive particle rearrangement, localized compaction, and the development of new contact networks throughout deformation (Fig. 10(a)). The color gradient indicates time, from the onset of failure (dark) to the final stage (light). This rate-strengthening evolution reflects the reorganization of the granular framework under increasing shear.
@@
Plotting friction against the inertial number $I_{n}$ shows that both dense and loose granular packings increase monotonically (Fig. 10(a)). The response of loose packings is more gradual, with a pronounced rise only at higher $I_{n}$. These patterns highlight the controlling role of packing density in strain localization, contact mechanics, and the frictional evolution of dry granular systems.
@@
While porosity in granular media is generally expected to vary with strain rate due to the competing effects of compaction and dilation, it can remain relatively stable over certain regimes. This constancy arises when the system achieves a balance between structural compression and dilation, or when strain rates are too low to significantly disrupt the grain arrangement. In the dense configuration, porosity remains nearly constant for inertial numbers below $10^{-1}$ (Fig. 10(c)). Parez2021 reported a similar trend, observing porosity stability across $I_{n}$ values ranging from $10^{-6}$ to $10^{-2}$, consistent with our findings. Likewise, Amarsid2024 found that porosity remains largely unaffected at low inertial numbers, further corroborating these observations for both dense and loose conditions.
V.2 Fluid-saturated condition
@@
Subaqueous granular rheology is set by the coupled interactions between solid grains and the surrounding fluid. In these systems, friction reflects both interparticle contact forces and hydrodynamic interactions, yielding greater complexity than in dry granular media. Loose subaqueous materials exhibit a strain-rate-dependent response with two regimes (Fig. 10(b)). At the onset of shearing, increasing strain rate drives pore-pressure buildup, which elevates friction via fluid drag and the rapid formation of reinforced contact networks. At lower strain rates—typically in the later stages of shearing—fluid lubrication dominates, facilitating particle rearrangement and reducing friction. The resulting rate-strengthening response underscores the central role of fluid-mediated forces in controlling deformation in loosely packed systems.
@@
Dense subaqueous packings respond differently: the friction coefficient decreases with increasing strain rate at shear onset (Fig. 10(b)), driven by dilation and the attendant negative pore pressure (suction). With further increases in rate, fluid drag becomes increasingly influential and pore-pressure effects intensify, reflecting the evolving coupling between the solid framework and the pore fluid. When friction is plotted against the viscous number $I_{v}$, dense packings approach a plateau at high $I_{v}$ with reduced variability, suggesting that viscous forces moderate the frictional response under steady fluid–particle interactions. In contrast, loose packings show a more pronounced rise in friction with increasing $I_{v}$, owing to intensified fluid drag and pore-pressure buildup.
@@
Although porosity in fluid-saturated granular materials is often expected to vary with strain rate, we observe that it remains nearly constant over a moderate range of rates (Fig. 10(d)). This invariance arises when compaction and dilation reach a dynamic balance, or when the imposed rates are too low to meaningfully perturb the microstructure. Consistent with these trends, Amarsid2024 reported that porosity—parameterized by the viscous number $I_{v}$—is largely insensitive at low $I_{v}$.
@@
A direct comparison of dry and saturated systems highlights the governing roles of fluid presence and initial fabric in the shear response: subaerial deformation is dominated by interparticle contact mechanics, whereas subaqueous deformation reflects coupled interactions among fluid drag, pore-pressure dynamics, and microstructural reorganization.
Building on these trends, we quantify the rheology by fitting data spanning the inertial and viscous regimes within the combined $K$-framework (Trulsson2012) (Fig. 11). The friction law is
$\mu(K)=\mu_{1}+\frac{(\mu_{2}-\mu_{1})\,K}{K+K_{0}},\quad\text{with }K=I_{v}+\alpha I_{n}^{2},$ (19)
where $\mu_{1}$, $\mu_{2}$, $K_{0}$, and $\alpha$ are fitting parameters.
Friction (Fig. 11) is well described by the rate-strengthening law in Eq. (19) across both subaerial and subaqueous regimes. To interpret this fit, we examine the relation between friction and packing fraction (Fig. 12) as a function of the inertial-to-viscous ratio, $x=\log_{10}(I_{n}/I_{v})$. The data delineate inertial- and viscous-dominated regimes in both environments, with a clear transition near $x=0$ ($I_{n}/I_{v}=1$). Plotting friction (Fig. 13(a)) and packing fraction (Fig. 13(b)) against $x$ cleanly separates the two regimes and reveals how the rheology evolves across subaerial and subaqueous conditions.
To directly compare the inertial and viscous contributions, Figs. 13(a,b) plot the friction and packing fraction against $x=\log_{10}(I_{n}/I_{v})$. For $x>0$ (inertial regime), $I_{n}$ dominates; for $x<0$ (viscous regime), $I_{v}$ dominates. In the inertial regime, friction increases while packing fraction decreases with shear rate, consistent with classical dry-granular rheology. In the viscous regime, packing fraction varies only weakly with rate, reflecting fluid-drag control and a diminished role for dilatancy. Together with Fig. 12 (friction versus packing fraction colored by $x$), these plots show that the material state spans a continuum between inertial- and viscous-dominated rheologies. This viewpoint naturally connects to the unified $K$-framework (Trulsson2012) (Fig. 11) and rationalizes the convergence of dense and loose flows when scaled by the dominant stress contribution. Overall, the frictional state is captured by the balance between inertial and viscous stresses, supporting the applicability of the unified rheological framework.
To further interrogate these trends, we compare our simulations and fitted parameters with those of Trulsson2012, who calibrated a similar rheology using a two-dimensional DEM of disks coupled to a continuum fluid in a simple–shear cell with periodic xx–boundaries, run at controlled normal stress $P_{p}$ and imposed mean shear velocity. By contrast, our analysis employs three-dimensional CFD–DEM column collapses of fluid-saturated beds with dense and loose initial packings, in which shear bands emerge naturally and stresses evolve transiently from grain motion rather than being servo–controlled. These differences in dimensionality, loading protocol, and stress control plausibly account for the deviations in our fitted $\mu(K)$ relative to simple–shear geometries.
Similarly, Trulsson2012 report simple-shear simulations in which the inertial-to-viscous number ratio spans $I_{n}/I_{v}\in[10^{-1},10^{4}]$ (i.e., $\log_{10}(I_{n}/I_{v})\in[-1,4]$). By contrast, our data span $I_{n}/I_{v}\in[10^{-6},10^{6}]$ (i.e., $\log_{10}(I_{n}/I_{v})\in[-6,6]$), extending the accessible range by seven decades (Fig. 12) and thereby capturing a broader spectrum of transient shear-zone dynamics during failure.
In the framework of Trulsson2012, viscous and inertial contributions are unified with a fitted coefficient $\alpha=0.635\pm 0.009$ and the constitutive form of Eq. (19), yielding $\mu_{1}=0.277\pm 0.001$, $\mu_{2}=0.85\pm 0.01$, and $\sqrt{K_{0}}=0.29\pm 0.01$. By contrast, fitting the same unified friction law (Eq. 19) to our data gives $\mu_{1}=0.378\pm 0.018$, $\mu_{2}=0.583\pm 0.055$, $\sqrt{K_{0}}=0.69\pm 0.59$, and $\alpha=1.232\pm 2.086$, with coefficients of determination $R^{2}=0.597$ (dense) and $R^{2}=0.814$ (loose). Both dense and loose cases are rate-strengthening, with loose systems attaining higher friction coefficients.
@@ VI Analytical solution for excess pore pressure using poromechanics
During breaching, a dynamic equilibrium emerges between pore pressure dissipation and continuous dilation You2012dynamics; Alhaddad2024-kj. As pore pressure dissipates, the effective stress within the slope deposit increases, stabilizing the sediment. However, for retrogressive failure to persist, this dissipation must be sustained by ongoing dilation. The resulting feedback mechanism, where slope failure induces dilative strain that regenerates negative excess pore pressure, helps maintain the breaching process. Importantly, this dynamic decouples the steady-state pore pressure distribution from the deposit’s intrinsic permeability, revealing an underexplored aspect of sediment failure behavior You2012dynamics; Alhaddad2024-kj.
To formalize this equilibrium, the analytical solution developed by You2012dynamics describes the pore pressure evolution coupled with dilation during the breaching process. This formulation predicts the steady-state negative pore pressure field, showing that the system can self-regulate pore pressure independently of permeability once dilation-driven feedback is active.
Comparisons with large-scale laboratory experiments and numerical models confirm this feedback mechanism’s role in sustaining breaching and highlight deviations during the transient initiation phase Alhaddad2024-kj; Alhaddad2023. The same analytical formulation is applied in this study to evaluate whether the CFD–DEM simulation results for excess pore pressure in both dense and loose subaqueous granular models follow the theoretical predictions, particularly at steady state. By extracting simulation data at various locations and times, we directly compare the evolution and final distribution of pore pressure against the analytical model.
Furthermore, recent hydromechanical models of subduction zone sediments Nikolinakou2023 demonstrate that shear-induced overpressures, in addition to changes in mean stress, significantly control pore pressure evolution. Although applied in a tectonic context, these findings conceptually reinforce that dilative strain and shear play a dominant role in transient and steady-state pore pressure fields, complementing observations from breaching processes.
@@
[…]
Eq. 20 describes pore pressure dissipation through Darcy flow with a moving boundary (first two terms) and continuous dilation effects (third term). The coefficient of consolidation, $C_{v}$, combines the effects of permeability, viscosity, and compressibility. $\hat{u}$ represents the excess pore pressure, $\hat{\sigma}_{3}$ is the least principal stress (assumed horizontal), and $C_{v}$ is the coefficient of consolidation. The boundary conditions applied in this analytical formulation are critical to reproducing the steady-state pore pressure field. In the Lagrangian reference frame, where the coordinate system moves with the retreating failure front, hydrostatic pore pressure is assumed at both the upstream (breaching front) and downstream (far-field) boundaries you2014mechanics. Additionally, no-flow conditions are imposed at the base and at the right boundary of the computational domain, assuming symmetry and minimal lateral drainage. These boundary constraints ensure that pore pressure evolves solely due to dilation-induced stress changes and fluid migration within the domain. The steady-state solution derived under these conditions shows that the excess pore pressure depends only on the dilation potential, effective stress, and retreat velocity, but not on permeability once equilibrium is reached. This insight reveals that boundary-controlled feedback mechanisms can dominate the evolution of excess pore pressure during breaching, consistent with both pore pressure measurements and model predictions you2014mechanics. Its solution, given by Eq. 21, represents $\hat{u}^{*}$, the excess pore pressure in the Lagrangian reference frame. The parameter $\beta$ denotes the relative dilation strength, defined as the ratio between differential stress-induced strain and mean effective stress-induced strain. $s_{0}$ represents the far-field horizontal stress. The coefficient of consolidation is defined as $C_{v}=\frac{k}{\mu E}$, where $k$ is the permeability, $\mu$ is the dynamic viscosity of the pore fluid, and $E$ is a modulus related to volumetric strain. The coordinate $\hat{x}$ is defined in the Lagrangian reference frame as $\hat{x}=x-vt$, where $v$ is the velocity of the retreating breaching front (Eq. 27). $\delta$ characterizes the ratio between $v$ and the product of the effective stress decay rate $\eta$ and the coefficient of consolidation $C_{v}$. The parameter $\eta$ is a constant with units of inverse length, affecting the spatial distribution of excess pore pressure.
@@
Figure 15: Evolution of excess pore pressure response during breaching and comparison with the analytical solution: (a) Monitored points A, B, D, E, and F for the subaqueous dense configuration; (d) monitored points G, H, J, K, and L for the subaqueous loose configuration. (b) In the dense configuration, wall removal results in an abrupt drop in pore pressure, followed by a gradual rise as the system stabilizes. (e) In contrast, the loose configuration exhibits an abrupt increase in pore pressure after wall removal, followed by a gradual decline. (c) and (f) Excess pore pressure at time intervals of 0.01 s, 0.1 s, 0.4 s, and 0.8 s (with dashed lines indicating the onset of breaching) is plotted against the distance from the breaching front using Lagrangian coordinates for the dense (point D) and loose (point J) configurations, respectively. The solid line represents the analytical solution. The comparison shows good agreement at steady state for both models: in the dense configuration, pore pressure begins at a negative value, increases, and stabilizes; in the loose configuration, pore pressure starts high, decreases, and reaches a stable value. Transient phases, however, show less agreement with the analytical model, highlighting the complexity of early-stage fluid–particle interactions.
@@
Eq. 23 defines permeability $k$ as a function of fluid properties and flow path geometry, where $v_{\text{avg}}$ is the average velocity, $L$ is the flow path length, and $\Delta P$ is the pressure difference. The calculated permeability values are $k=1.75\times 10^{-7}\,\mathrm{m}^{2}$ for the loose configuration and $k=0.25\times 10^{-7}\,\mathrm{m}^{2}$ for the dense configuration. The evolution of permeability over time at selected points, derived using Darcy’s law, is presented in Supplementary Fig. S1(a). In Eq. 24, $\beta$ is the dimensionless dilation strength, where $\beta>\frac{1}{2}$ indicates dilative material. Eq. 26 describes the least principal stress, which decays exponentially with distance from the breaching front, explaining localized pore pressure changes during continuous dilation. During breaching, the volumetric strain rate equals the water flux into the deposit, and strain is proportional to the minimum pore pressure. The parameter $s_{0}$ represents the least principal stress as $x\to\infty$, and $\eta$ defines the rate of stress decay.